# Bounds for Lyapunov exponent of circular light orbits in black holes

### v2

@@ Abstract
Chaotic systems near black holes satisfy a universal bound, $\lambda\leq\kappa_{H}$ linking the Lyapunov coefficient $\lambda$ associated with unstable orbits to surface gravity $\kappa_{H}$ of the event horizon. A natural question is whether this bound is satisfied by unstable circular null geodesics in the vicinity of black holes. However, there are known cases where this bound is violated. It is intriguing to ask whether there exists an alternative universal bound that is valid in such situations. We show that for any spherically symmetric, static black hole that satisfies Einstein’s equations and the dominant energy condition, there exist other universal bounds relating the Lyapunov coefficient to a generalized notion of surface gravity at the photon sphere, as well as to the Unruh temperature locally measured by static observers. As applications, we show how these bounds also constrain the imaginary part of quasinormal modes in the eikonal regime and how the Lyapunov coefficient relates to the shadow size and the entropy of the horizon.
@@ I Introduction
Nevertheless, Abreu and Visser (2010) introduced a concept of a generalized surface gravity, $\kappa(r)$, for static, spherically symmetric spacetimes, which is applicable at arbitrary radial positions, $r$. In particular, evaluation of $\kappa(r)$ at the position $r_{H}$ of a black hole horizon yields its surface gravity $\kappa(r_{H})=\kappa_{H}$. Moreover, at the radial position $r_{\gamma}$ of unstable circular null geodesics, it is non-vanishing even for nearly extremal black holes. This provokes the question if a similar relation to eq. (1) holds when replacing $\kappa_{H}$ by the generalized surface gravity $\kappa_{\gamma}\equiv\kappa(r_{\gamma})$. That is, setting a bound to the Lyapunov exponent of circular null geodesics with the value of the corresponding generalized surface gravity at $r_{\gamma}$.
@@
Our answer to this question is affirmative. For any spherically symmetric black hole satisfying the Einstein equations and the dominant energy condition, there exists a universal upper bound relating $\lambda$ and $\kappa_{\gamma}$. This upper bound also explicitly depends on the radial position of the photon sphere $r_{\gamma}$. Furthermore, we show the existence of a bound between the Lyapunov exponent and the Unruh temperature $T_{U}$ measured by static observers located at $r_{\gamma}$. This bounding relation, $\lambda\leq 2\pi T_{U}$, is saturated for a Schwarzschild black hole. In fact, it was recently derived in Giataganas et al. (2024) for the two particular cases of Schwarzschild and Reissner-Nordström black holes. Our proof generalizes their results to a broader class of spherically symmetric black holes and is sufficiently general to not require an explicit form of the metric. Moreover, our new bound is directly related to recent bounds associated with the shadow size and acceleration bounds for radial linear uniformly accelerated trajectories in static spherically symmetric black hole spacetimes of the Schwarzschild type Paithankar and Kolekar (2023).
@@
Our major results culminate in the expressions for the generic bounds between the Lyapunov exponent and (i) surface gravity at the photons sphere (eq. (38), (39) and (40)), (ii) the Unruh temperature of static observers at the photon sphere (eq. (42)), and (iii) the shadow of the black hole (eq. (44)).
@@ III Bounds on the Lyapunov exponent of circular null geodesics
To establish a bound relating the Lyapunov exponent with a surface gravity, which does not exclude these cases, we employ the notion of the generalized surface gravity $\kappa(r)$ of Abreu and Visser (2010) which is valid for every sphere of radius $r$. The usage of $\kappa(r)$ provides simple bounds on quasi-local entropies similar to the holographic bounds relating entropy with area of the enclosed region. This generalized surface gravity $\kappa(r)$ is determined by a redshift factor of the norm $a=(g_{bc}a^{b}a^{c})^{1/2}$ of the acceleration $a^{b}=t^{c}\nabla_{c}t^{b}$ of observers following the integral curves along the unit vector field $t^{c}$. For (2), it results in
@@
where $a_{U}=a(r_{\gamma})$. This norm of the acceleration at $r_{\gamma}$ relates to the Unruh temperature $T_{U}(r_{\gamma})$, which is minimum local temperature of vacuum measured by static observers at $r_{\gamma}$
@@
It turns out that the relations (38), (39) and (40) are not the only inequalities which can be written at the photon sphere of spherically symmetric black hole under the given physical assumption. In fact, we can re-express the inequality (38) in a very natural way in terms of the Unruh temperature as measured by static observers placed at $r_{\gamma}$. To do so, note that from eq. (34) and the dominant energy condition, it follows that
@@
$\lambda\leq 2\pi T_{U}.$ (42)
It is worth noting that, recently, in reference Giataganas et al. (2024), the validity of this inequality was observed for the special cases of Schwarzschild and Reissner-Nordström black holes through direct calculation. What we have shown in Eq.(42) is that it remains valid within Einstein’s theory of general relativity, under very natural physical assumptions. In Giataganas et al. (2024), the authors also conjecture new inequalities that have been tested to hold for rotating spacetimes in the Kerr-Newman family (more on this will be discussed below). Note also, that a relation between the Lyapunov exponent and the Unruh temperature associated to a scalar field thermalization in the vicinity of an effective metric near to the photon sphere was also observed in Raffaelli (2022).
@@
where $S$ is obtained from the area $A$ of the event horizon via $S=\frac{A}{4}=\pi r_{H}^{2}$.
@@
which is a weaker bound than (47). The inequality (48) holds for a broader family of spherically symmetric black holes than (47) of Giataganas et al. (2024). Their inequality was shown to remain valid for the Kerr-Newman family of black holes. Since (48) is a weaker bound, it is clear that it remains valid for the spacetimes studied in Giataganas et al. (2024). Naturally, for spherically symmetric black holes, writing the upper bound for $\lambda$ in terms of $r_{H}$ or $S$ is entirely equivalent. However, this equivalence does not hold for spinning black holes. It would be intriguing to establish a general proof of the validity of some form of these inequalities for rotating black holes that satisfy the Einstein equations and the dominant energy condition. Alternatively, it would be valuable to identify counterexamples where one or the other of these proposed inequalities is violated.
@@ IV Summary
We have proven that for every static and spherically symmetric black hole satisfying the Einstein equations and the dominant energy condition, there exists a set of universal bounds for the Lyapunov exponent $\lambda$ associated to unstable circular null geodesics in terms of the generalized surface gravity $\kappa_{\gamma}$ Eq.(39) and (40); the Unruh temperature measured by static observers at $r_{\gamma}$, Eq.(42); and the shadow size $R_{sh}$ Eq.(44). Furthermore, we have reformulated some of these bounds in terms of the black hole’s entropy Eq.(48).
@@
It would be interesting to know whether some of the bounds proven here have analogs in other spherically symmetric black holes that do not satisfy the assumptions made in this work, specifically regarding the energy conditions or the requirement of being solutions to the Einstein equations. For instance, a generalized version of (19) was proven in for the $n$-dimensional Einstein-Gauss-Bonnet theory (see also Ghosh et al. (2023)), and in Cvetic et al. (2016) Cvetic, Gibbons and Pope have shown that black holes arising from other gravitational theories also satisfy the bounds found inGallo and Villanueva (2015). Since the Lyapunov bound is, in some sense, related to the upper bound on the location of the photon sphere, it is hoped that more general versions of the bound for $\lambda$ exist for broader classes of metrics or alternative theories. More importantly, it would be desirable to test these inequalities for rotating black holes or to prove their validity in those spacetimes. However, in general, the Einstein equations for rotating spacetimes are more complex, and the null geodesics are typically non-planar. Therefore, further work and potentially different techniques will be required to address this problem.

### v3

@@ Abstract
Chaotic systems near black holes satisfy a universal bound, $\lambda\leq\kappa_{H}$ linking the Lyapunov coefficient $\lambda$ associated with unstable orbits to surface gravity $\kappa_{H}$ of the event horizon. A natural question is whether this bound is satisfied by unstable circular null geodesics in the vicinity of black holes. However, there are known cases where this bound is violated. It is intriguing to ask whether there exists an alternative universal bound that is valid in such situations. We show that for any spherically symmetric, static black hole that satisfies Einstein’s equations and the dominant energy condition, there exist other universal bounds relating the Lyapunov coefficient to a generalized notion of surface gravity at the photon sphere. As applications, we show how these bounds also constrain the imaginary part of quasinormal modes in the eikonal regime and how the Lyapunov coefficient relates to the shadow size and the entropy of the horizon.
@@ I Introduction
Nevertheless, Abreu and Visser (2010) introduced a concept of a “generalized surface gravity”, $\kappa(r)$, for static, spherically symmetric spacetimes, which is applicable at arbitrary radial positions111 In Abreu and Visser (2010), the concept of generalized surface gravity was introduced in the context of a general static spacetime. For our purposes, we limit their study to static and spherically symmetric black holes. $r$. In particular, evaluation of $\kappa(r)$ at the position $r_{H}$ of a black hole horizon yields its surface gravity $\kappa(r_{H})=\kappa_{H}$. As shown in Abreu and Visser (2010) the usage of $\kappa(r)$ provides simple bounds on quasi-local entropies similar to the holographic bounds relating the entropy $S(r)$ with area $A(r)$ of the enclosed region, namely,
$S(r)\leq\frac{\kappa(r)}{\kappa_{H}}\frac{A(r)}{2},$ (2)
with $\kappa_{H}$ the standard surface gravity associated to the event horizon.
@@
Moreover, at the radial position $r_{\gamma}$ of unstable circular null geodesics, $\kappa(r_{\gamma})$ is non-vanishing even for nearly extremal black holes. This provokes the question if a similar relation to eq. (1) holds when replacing $\kappa_{H}$ by the generalized surface gravity $\kappa_{\gamma}\equiv\kappa(r_{\gamma})$. That is, setting a bound to the Lyapunov exponent of circular null geodesics with the value of the corresponding generalized surface gravity at $r_{\gamma}$.
@@
Our answer to this question is affirmative. For any spherically symmetric black hole satisfying the Einstein equations and the dominant energy condition, there exists a universal upper bound relating $\lambda$ and $\kappa_{\gamma}$. This upper bound also explicitly depends on the radial position of the photon sphere $r_{\gamma}$. Furthermore, we show the existence of a bound between the Lyapunov exponent and an operational Unruh temperature $T_{U}=\frac{a_{U}}{2\pi}$ with $a_{U}$ being the norm of the proper acceleration of static observers located at $r_{\gamma}$ We refer to this temperature as operational because, in reality, a static observer near a black hole would measure a different temperature, namely the Hawking radiation temperature adjusted by a redshift factor that depends on the observer’s location. However, this operational temperature emerges in effective geometries associated with the photon sphere Raffaelli (2022); Giataganas et al. (2024). To facilitate comparison with those works we have chosen to express some of the derived inequalities in terms of $T_{U}$. Notably, in Barbado et al. (2016), the authors demonstrated that the temperature $T_{U}$ is indeed measured by radially free-falling observers starting from rest at a radius $r_{\gamma}$ for radiation in the so-called ingoing sector [^2] While Barbado et al. (2016) computations are valid for any fixed radius $r$, we specifically refer to $r_{\gamma}$ as our discussion is confined to the photon sphere region. (i.e., radiation pointing towards the black hole). That is, while a static observer with acceleration $a_{U}$ detects Hawking radiation, a radially free-falling observer instantaneously at rest there measures ingoing radiation with a temperature $T_{U}$ (at the initial time, when the observer is at rest [^3] These observers also measures outgoing radiation with contribution of Hawking and Unruh effect. For other positions of the radially free-falling observer, the temperature of the ingoing radiation has Doppler corrections due to the relative velocity between the free-falling observer and static observers (see Barbado et al. (2016) and references therein for more details).), which precisely matches $\frac{a_{U}}{2\pi}$. The bounding relation between Lyapunov exponent and operational Unruh temperute, $\lambda\leq 2\pi T_{U}$, is saturated for a Schwarzschild black hole. In fact, it was recently derived in Giataganas et al. (2024) for the two particular cases of Schwarzschild and Reissner-Nordström black holes. Our proof generalizes their results to a broader class of spherically symmetric black holes and is sufficiently general to not require an explicit form of the metric. Moreover, our new bound is directly related to recent bounds associated with the shadow size and acceleration bounds for radial linear uniformly accelerated trajectories in static spherically symmetric black hole spacetimes of the Schwarzschild type Paithankar and Kolekar (2023).
@@
Our major results culminate in the expressions for the generic bounds between the Lyapunov exponent and (i) surface gravity at the photons sphere (eq. (39), (40) and (41)), (ii) the acceleration of static observers at the photon sphere (eq. (43)), and (iii) the shadow of the black hole (eq. (45)).
@@ III Bounds on the Lyapunov exponent of circular null geodesics
To establish a bound relating the Lyapunov exponent with a surface gravity, which does not exclude these cases, we employ the notion of the generalized surface gravity $\kappa(r)$ of Abreu and Visser (2010) which is valid for every sphere of radius r. It is important to clarify that this concept of surface gravity does not imply the existence of a quasilocal version of the first law of thermodynamics at fixed radii $r_{o}$, relating changes in energy $\delta E$ of infalling objects crossing the timelike surface $r=r_{o}$ (as measured by static observers at that location) to changes in the area $\delta A$ of the black hole’s event horizon through $\delta E=\frac{{\kappa(r_{o})}}{8\pi}\delta A$. Although a local version of this law exists in the literature Frodden et al. (2013); Tripathy et al. (2024), it refers to an effective surface gravity given by $\tilde{\kappa}=\frac{\kappa_{H}}{|t^{a}|_{r_{o}}}$, with $|t^{a}|_{r_{o}}$ the norm of the timelike Killing vector $t^{a}$ at $r=r_{o}$. However, as shown in Ravuri and McMaken (2025), $\kappa(r)$ contributes to local temperatures of radially free-falling observers. A generalization of $\kappa(r)$ for non-static spacetimes can be found in Barcelo et al. (2011). This generalized surface gravity $\kappa(r)$ is determined by a redshift factor of the norm $a=(g_{bc}a^{b}a^{c})^{1/2}$ of the acceleration $a^{b}=t^{c}\nabla_{c}t^{b}$ of observers following the integral curves along the unit vector field $t^{c}$. For (3), it results in
@@
where $a_{U}=a(r_{\gamma})$. This norm of the acceleration at $r_{\gamma}$ relates to $T_{U}(r_{\gamma})$ by
@@
It turns out that the relations (39), (40) and (41) are not the only inequalities which can be written at the photon sphere of spherically symmetric black hole under the given physical assumption. In fact, we can re-express the inequality (39) in a very natural way in terms of the “Unruh temperature” $T_{U}$. To do so, note that from eq. (35) and the dominant energy condition, it follows that
@@
$\lambda\leq a_{U}=2\pi T_{U}.$ (43)
It is worth noting that, recently, in reference Giataganas et al. (2024), the validity of this inequality was observed for the special cases of Schwarzschild and Reissner-Nordström black holes through direct calculation. What we have shown in Eq.(43) is that it remains valid within Einstein’s theory of general relativity, under very natural physical assumptions. Note also, that a relation between the Lyapunov exponent and the Unruh temperature associated to a scalar field thermalization in the vicinity of an effective metric near to the photon sphere was also observed in Raffaelli (2022).
@@
The authors’ motivation for this empirical formula was to establish a new inequality that remains valid for a larger family of black holes. In particular, they observed that the inequality $\lambda\leq\kappa_{H}$ which holds for the Schwarzschild spacetime, can be rewritten as (48) by noting that in the Schwarzschild case, the surface gravity is given by $\kappa_{H}=\frac{1}{2r_{H}}$, and the entropy is expressed as
$S=\frac{A(r_{H})}{4}=\pi r_{H}^{2}.$ (49)
This allowed them to explore its validity in other spacetimes. As a result, they showed that (48) holds for the whole Kerr-Newman family of black holes.
@@
which is a weaker bound than (48). The inequality (50) holds for a broader family of spherically symmetric black holes than (48) of Giataganas et al. (2024). Since (50) is a weaker bound, it is clear that it remains valid for the spacetimes studied in Giataganas et al. (2024). Naturally, for spherically symmetric black holes, writing the upper bound for $\lambda$ in terms of $r_{H}$ or $S$ is entirely equivalent. However, this equivalence does not hold for spinning black holes. It would be intriguing to establish a general proof of the validity of some form of these inequalities for rotating black holes that satisfy the Einstein equations and the dominant energy condition. Alternatively, it would be valuable to identify counterexamples where one or the other of these proposed inequalities is violated.
@@ IV Summary
We have proven that for every static and spherically symmetric black hole satisfying the Einstein equations and the dominant energy condition, there exists a set of universal bounds for the Lyapunov exponent $\lambda$ associated to unstable circular null geodesics in terms of the generalized surface gravity $\kappa_{\gamma}$ Eq.(40) and (41); the Unruh temperature Eq.(43); and the shadow size $R_{sh}$ Eq.(45). Furthermore, we have reformulated some of these bounds in terms of the black hole’s entropy Eq.(50). However, to our knowledge, a justification for the validity (or invalidity) of inequalities like (50) based on fundamental physical principles remains unknown. Exploring possible underlying reasons could be an interesting direction for future research.
@@
It would be interesting to know whether some of the bounds proven here have analogs in other spherically symmetric black holes that do not satisfy the assumptions made in this work, specifically regarding the energy conditions or the requirement of being solutions to the Einstein equations. For instance, a generalized version of (20) was proven in Gallo and Villanueva (2015) for the $n$-dimensional Einstein-Gauss-Bonnet theory (see also Ghosh et al. (2023)), and in Cvetic et al. (2016) Cvetic, Gibbons and Pope have shown that black holes arising from other gravitational theories also satisfy the bounds found inGallo and Villanueva (2015). Since the Lyapunov bound is, in some sense, related to the upper bound on the location of the photon sphere, it is hoped that more general versions of the bound for $\lambda$ exist for broader classes of metrics or alternative theories. More importantly, it would be desirable to test these inequalities for rotating black holes or to prove their validity in those spacetimes. However, in general, the Einstein equations for rotating spacetimes are more complex, and the null geodesics are typically non-planar. Therefore, further work and potentially different techniques will be required to address this problem.